OpenStudy (anonymous):

What are the solutions of the quadratic equation below? 3x^2 - x = 11

OpenStudy (anonymous):

A:\[1\pm \sqrt{133\div6}\] B:\[-1\pm \sqrt{133\div6}\] C:\[\pm \sqrt{133\div6}\] D.\[\pm \sqrt{133\div3}\]

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

Why don't you plug it into the quadratic formula?

OpenStudy (anonymous):

i dont know the formuls

OpenStudy (mertsj):

a=3 b=-1 c=-11

OpenStudy (mertsj):

\[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (mertsj):

That is the formula